Malma is a HK manufacturer that has recently changed suppliers and has started importing one of its key components from Germany. The Finance Director has suggested that the company hedge all future Euro foreign currency transactions however the Managing Director thinks that the Euro is very stable and that this is a cost that could be avoided.

The first shipment has arrived and Malma are due to pay 2,700,000 Euros in 3 months’ time. The following information has been provided in order to assess the use of a forward contract and a money market hedge:

                                           

Current spot exchange rate       0.1285 – 0.1290 Euro / HK$

3 month future exchange rate    0.1288 – 0.1294 Euro / HK$

Annual Euro borrowing rate       6% per annum

Annual Euro deposit rate            3% per annum

Annual HK deposit rate               4% per annum

Annual HK borrowing rate          7% per annum

Malma currently has a cash surplus.

Required:

  1. Calculate the HK$ payment for Malma using a forward contract & money market hedge.
Answer #1

HK$ Payment using a forward contract:

3 month future exchange rate =(0.1288+0.1294)/2=0.1291Euro/HK$

Amount of Euro to be paid=2,700,000

Amount to be paid in HK$=2700000/0.1291=20,914,020

HK$ payment using money market hedge:

Current Spot Rate =(0.1285+0.1290)/2=0.12875 Euro/HK$

Euro deposit rate=3% per annum

Deposit rate for 3 months=3/4=0.75%

Present Value of 2700000Euro to be paid after 3 months=2700000/(1+0.0075)=2,679,901 Euro

Amount of HK$ at the current spot rate =2679901/0.12875=20,814,763 HK$

HK borrowing rate for 3 months =(7/4)%=1.75%

Future Value of HK $ 20,814,763 borrowed =20814763*(1+0.0175)=21,179,021

HK$ Payment using money market hedge =21,179,021

Strategy:

Borrow HK $ 20,814,763

Convert at Current spot rate to 2,679,901 Euro

Deposit at 3% rate in Euro

Future value of deposit =2700000 Euro

Amount to be paid on borrowings in HK$=21,179,021